# Acylindrical hyperbolicity and the centers of Artin groups that are not free of infinity

### v2

@@ 2.2. Defining graphs of Artin groups
We now have a finite simple labeled graph $\Gamma$ with the vertex set $V(\Gamma)=V$, the edge set $E(\Gamma)=\bigsqcup_{m\in\mathbb{Z}_{\geq}2}\tilde{\mu}^{-1}(m)$, and the labeling $\mu:=\tilde{\mu}|_{E(\Gamma)}$. The Artin group $A_{\Gamma}$ associated with $\Gamma$ is then defined by the presentation (1.1) and $\Gamma$ is called the defining graph of $A_{\Gamma}$.
For convenience, we define two other finite simple graphs $\Gamma^{c}$ and $\Gamma^{t}$ as follows. $\Gamma^{c}$ is the finite simple graph with the vertex set $V(\Gamma^{c})=V$ and the edge set $E(\Gamma^{c})=\tilde{\mu}^{-1}(\infty)$. This is the so-called complement graph of $\Gamma$. $\Gamma^{t}$, the Coxeter graph, is the finite simple graph with the vertex set $V(\Gamma^{t})=V$ and the edge set $E(\Gamma^{t})=\bigsqcup_{m\in\mathbb{Z}_{\geq 3}\cup\{\infty\}}\tilde{\mu}^{-1}(m)$.
@@ 2.3. Properties of Artin groups
We collect and show some properties of Artin groups. Let $A_{\Gamma}$ be an Artin group associated with a defining graph $\Gamma$ with the vertex set $V(\Gamma)=V$ and the edge set $E(\Gamma)=E$, as in Section 2.2. Lemma 2.11 is essentially used in the proof of Theorem 1.2.
By study of van der Lek (see [35], [41]), the following are well-known, which often used without mention in this paper.
@@
    For any subset $U\subset V$, the subgroup of $A_{\Gamma}$ generated by $U$, called a standard parabolic subgroup, is itself an Artin group associated with the full subgraph of $\Gamma$ spanned by $U$. (We denote this subgroup as $A_{U}$. When $U$ is empty, we define $A_{\emptyset}=\{1\}$.)
@@
Take an element $v\in V$ and a subset $V^{\prime}\subset V$. Suppose that $v\not\in V^{\prime}$. Set $V^{\prime\prime}:=\{v^{\prime\prime}\in V^{\prime}\mid\mu(v,v^{\prime\prime})=2\}$. Then $A_{V^{\prime}}\cap vA_{V^{\prime}}v^{-1}=A_{V^{\prime\prime}}$. In particular for any $v^{\prime}\in V^{\prime}\setminus V^{\prime\prime}$, $A_{V^{\prime}}\cap vA_{V^{\prime}}v^{-1}\subset A_{V^{\prime}\setminus\{v^{\prime}\}}$.
@@
Clearly we have $A_{V^{\prime\prime}}\subset A_{V^{\prime}}\cap vA_{V^{\prime}}v^{-1}$. By noting that $v$ is $(V^{\prime},V^{\prime})$-reduced in the Coxeter group $W_{\Gamma}$ and applying [19, Proposition 3.3] to our setting, we have some $V^{\prime\prime\prime}\subset V^{\prime}$ such that $vA_{V^{\prime}}v^{-1}\cap A_{V^{\prime}}=A_{V^{\prime\prime\prime}}$. Hence $V^{\prime\prime}\subset V^{\prime\prime\prime}$. Assume that $V^{\prime\prime}\subsetneq V^{\prime\prime\prime}$. Take $v^{\prime}\in V^{\prime\prime\prime}\setminus V^{\prime\prime}$. Then $v^{\prime}\in A_{V^{\prime\prime\prime}}=vA_{V^{\prime}}v^{-1}\cap A_{V^{\prime}}$ and thus $v^{-1}v^{\prime}v\in A_{V^{\prime}}$. By $v^{-1}v^{\prime}v\in A_{\{v,v^{\prime}\}}$ and $\{v,v^{\prime}\}\cap V^{\prime}=\{v^{\prime}\}$, we have $v^{-1}v^{\prime}v\in A_{\{v,v^{\prime}\}}\cap A_{V^{\prime}}=A_{\{v^{\prime}\}}$. Therefore we have $m\in\mathbb{Z}$ such that $v^{-1}v^{\prime}v=(v^{\prime})^{m}$. Since the surjection $A_{\{v,v^{\prime}\}}\to A_{\{v^{\prime}\}}$ defined by $v,v^{\prime}\mapsto v^{\prime}$ gives $v^{-1}v^{\prime}v\mapsto v^{\prime}$ and $(v^{\prime})^{m}\mapsto(v^{\prime})^{m}$, we see $m=1$. Hence we have $v^{-1}v^{\prime}v=v^{\prime}$ in $A_{\Gamma}$. This means $\mu(v,v^{\prime})=2$. This contradicts $v^{\prime}\notin V^{\prime\prime}$. Hence $V^{\prime\prime\prime}=V^{\prime\prime}$. Therefore $vA_{V^{\prime}}v^{-1}\cap A_{V^{\prime}}=A_{V^{\prime\prime}}$. ∎
@@ 2.4. Clique-cube complexes and actions on them
The clique-cube complex $C_{\Gamma}$ is the cube complex whose vertices (i.e., 00-dimensional cubes) are cosets $gA_{U}$ ($g\in A_{\Gamma},U\in\Delta_{\Gamma}$), where two vertices $gA_{U}$ and $hA_{U^{\prime}}$ are joined by an edge (i.e., a $1$-dimensional cube) in $C_{\Gamma}$ if and only if $gA_{U}\subset hA_{U^{\prime}}$ and $U$ and $U^{\prime}$ differ by a single generator. Note that, in this case, we can always replace $h$ by $g$, that is, $hA_{U^{\prime}}=gA_{U^{\prime}}$. More generally, two vertices $gA_{U}$ and $gA_{U^{\prime}}$ with $gA_{U}\subset gA_{U^{\prime}}$ span a #$\#(U^{\prime}\setminus U)$-dimensional cube $[gA_{U},gA_{U^{\prime}}]$ in $C_{\Gamma}$.
The group $A_{\Gamma}$ acts on the clique-cube complex $C_{\Gamma}$ by left multiplication, $h\cdot gA_{U}=(hg)A_{U}$. This action preserves the cubical structure and is isometric. The action is also co-compact with a fundamental domain $\bigcup_{U\in\Delta_{\Gamma}}[A_{\emptyset},A_{U}]$, where $[A_{\emptyset},A_{U}]$ is a #$\#U$-dimensional cube spanned by two vertices $A_{\emptyset}$ and $A_{U}$ in $C_{\Gamma}$. However, the action is not proper. In fact, the stabilizer of a vertex $gA_{U}$ is the conjugate subgroup $gA_{U}g^{-1}$, so all vertices except translations of $A_{\emptyset}$ have infinite stabilizers. We also note that $C_{\Gamma}$ is not a proper metric space because it contains infinite valence vertices. Additionally, $C_{\Gamma}$ has infinite diameter if and only if $\Gamma$ itself is not a clique.
@@
Each edge in $C_{\Gamma}$ can be labeled with a generator in $V$. For example, the edge between $gA_{U}$ and $gA_{U\sqcup\{v\}}$ is labeled by $v$. Any two parallel edges in a cube have the same label, so we can also label the hyperplane dual to such an edge by $v$ and say that such a hyperplane is of $v$-type. Every hyperplane of $v$-type is the translation of a hyperplane dual to the edge between $A_{\emptyset}$ and $A_{\{v\}}$. If a hyperplane of $v$-type crosses another hyperplane of $v^{\prime}$-type, then $(v,v^{\prime})\in E(\Gamma)$. In particular, two different hyperplanes of the same type do not cross each other.
We endow a metric on $C_{\Gamma}$ in the usual way as a cube complex. That is, we identify all cubes in $C_{\Gamma}$ with Euclidean unit cubes and give $C_{\Gamma}$ the path-metric.
Theorem 2.14 ([14, Theorem 2.2]).
The clique-cube complex $C_{\Gamma}$ is a complete CAT(0) space.
@@ 2.5. Joins of graphs
A simple graph $\Gamma$ is said to be decomposable as a join if there exist an index set $\Lambda$ with #$\#\Lambda\geq 2$ and subgraphs $\Gamma_{\alpha}$ ($\alpha\in\Lambda$) of $\Gamma$ such that $\Gamma=\ast_{\alpha\in\Lambda}\Gamma_{\alpha}$. This is called a join decomposition of $\Gamma$ into factors $\Gamma_{\alpha}$ ($\alpha\in\Lambda$). $\Gamma$ is said to be indecomposable as a join if it is not decomposable as a join.
@@
Any simple graph $\Gamma$ is indecomposable as a join if and only if its complement graph $\Gamma^{c}$ is connected.
@@
Let $\Gamma$ be a simple graph. Suppose that $\Gamma$ is decomposable as a join. For the unique join decomposition into indecomposable factors $\Gamma=\ast_{\alpha\in\Lambda}\Gamma_{\alpha}$, we set $\Lambda_{0}:=\{\alpha\in\Lambda\mid\ |\Gamma_{\alpha}|=1\}$ and $\Gamma_{0}:=\ast_{\alpha\in\Lambda_{0}}\Gamma_{\alpha}$. We call $\Gamma_{0}$ the maximal clique factor of $\Gamma$. Also we set $\Lambda_{\ast}:=\Lambda\setminus\Lambda_{0}$ and $\Gamma_{\ast}:=\ast_{\alpha\in\Lambda_{\ast}}\Gamma_{\alpha}$. We call $\Gamma$ a cone if $\Lambda_{0}\neq\emptyset$.
See Figures 1 and 2. We have $\Gamma=\Gamma_{0}\ast\Gamma_{\ast}$. Clearly $\Gamma$ is a clique if and only if $\Lambda=\Lambda_{0}$.
@@ 3. Reduced clique-cube complexes and actions on them
In this section, we introduce reduced clique-cube complexes and study them. Let $A_{\Gamma}$ be an Artin group associated with a defining graph $\Gamma$ with the vertex set $V(\Gamma)=V$ and the edge set $E(\Gamma)=E$, as in Section 2.2. We consider the unique join decomposition $\Gamma=\Gamma_{0}\ast\Gamma_{\ast}=\Gamma_{0}\ast\left(\ast_{1\leq i\leq k}\Gamma_{i}\right)$ where $\Gamma_{0}$ is the maximal clique factor and $\Gamma_{i}$ $(1\leq i\leq k)$ is an indecomposable factor with at least two vertices (see Definition 2.18). We set $V_{i}=V(\Gamma_{i})$.
@@
$\Delta_{\Gamma,\Gamma_{0}}=\{U\subset V\mid U\in\Delta_{\Gamma},\ U\supset V(\Gamma_{0})\},$
where $\Gamma_{0}$ is the maximal clique factor of $\Gamma$. Let $C_{\Gamma,\Gamma_{0}}$ be a subcomplex of $C_{\Gamma}$ spanned by vertices $gA_{U}$ with $g\in A_{\Gamma}$ and $U\in\Delta_{\Gamma,\Gamma_{0}}$.
When $\Gamma$ is not a cone, that is, $\Gamma=\Gamma_{\ast}$, we have $C_{\Gamma,\Gamma_{0}}=C_{\Gamma}$. When $\Gamma$ is a clique, that is, $\Gamma=\Gamma_{0}$, we have $C_{\Gamma,\Gamma_{0}}=C_{\Gamma,\Gamma}=\{A_{\Gamma}\}$. In this paper, we treat the case where $\Gamma$ is not a clique.
We note that the action of $A_{\Gamma}$ on $C_{\Gamma}$ preserves $C_{\Gamma,\Gamma_{0}}$. In the following, we consider the induced action of $A_{\Gamma}$ on $C_{\Gamma,\Gamma_{0}}$.
@@
If $\Gamma$ is a clique, then $C_{\Gamma,\Gamma_{0}}=\{A_{\Gamma}\}$ is clearly connected. Suppose that $\Gamma$ is not a clique. Let $U\in\Delta_{\Gamma,\Gamma_{0}}$. Let $g\in A_{\Gamma}$. First, we show that there exists a path in $C_{\Gamma,\Gamma_{0}}$ from $A_{V(\Gamma_{0})}$ to $gA_{V(\Gamma_{0})}$. Suppose that $g\notin A_{V(\Gamma_{0})}$, since otherwise $A_{V(\Gamma_{0})}=gA_{V(\Gamma_{0})}$. Let $g=v_{1}^{\varepsilon_{1}}\cdots v_{n}^{\varepsilon_{n}}$ where $v_{i}\in V(\Gamma)$, $\varepsilon_{i}\in\{1,-1\}$ for every $1\leq i\leq n$. By induction on $n$, we assume that there exists a path in $C_{\Gamma,\Gamma_{0}}$ from $A_{V(\Gamma_{0})}$ to $v_{1}^{\varepsilon_{1}}\cdots v_{n-1}^{\varepsilon_{n-1}}A_{V(\Gamma_{0})}$. If $v_{n}\in V(\Gamma_{0})$, then
$v_{1}^{\varepsilon_{1}}\cdots v_{n-1}^{\varepsilon_{n-1}}A_{V(\Gamma_{0})}=v_{1}^{\varepsilon_{1}}\cdots v_{n}^{\varepsilon_{n}}A_{V(\Gamma_{0})},$
and the induction proceeds. If $v_{n}\notin V(\Gamma_{0})$, since $V(\Gamma_{0})\cup\{v_{n}\}\in\Delta_{\Gamma,\Gamma_{0}}$, there exists a path
$\left(v_{1}^{\varepsilon_{1}}\cdots v_{n-1}^{\varepsilon_{n-1}}A_{V(\Gamma_{0})},\ v_{1}^{\varepsilon_{1}}\cdots v_{n-1}^{\varepsilon_{n-1}}A_{V(\Gamma_{0})\cup\{v_{n}\}}=gA_{V(\Gamma_{0})\cup\{v_{n}\}},\ gA_{V(\Gamma_{0})}\right)$
from $A_{V(\Gamma_{0})}$ to $gA_{V(\Gamma_{0})}$ in $C_{\Gamma,\Gamma_{0}}$. Next, we show that there exists a path in $C_{\Gamma,\Gamma_{0}}$ from $gA_{V(\Gamma_{0})}$ to $gA_{U}$. Let $U=V(\Gamma_{0})\cup\{u_{1},\ldots,u_{m}\}$ where $u_{i}\neq u_{j}$ for $1\leq i\neq j\leq m$. Then there exists a path
$\left(gA_{V(\Gamma_{0})},gA_{V(\Gamma_{0})\cup\{u_{1}\}},gA_{V(\Gamma_{0})\cup\{u_{1},u_{2}\}},\ldots,gA_{U}\right)$
in $C_{\Gamma,\Gamma_{0}}$ from $gA_{V(\Gamma_{0})}$ to $gA_{U}$. Therefore, it follows that $C_{\Gamma,\Gamma_{0}}$ is a connected subcomplex of $C_{\Gamma}$. ∎
@@
Let $X$ be a geodesic space and $W$ be a subset of $X$. $W$ is convex in $X$ if for every $w_{1},w_{2}\in W$, any geodesic from $w_{1}$ to $w_{2}$ is contained in $W$.
@@
$C_{\Gamma,\Gamma_{0}}$ is convex in $C_{\Gamma}$. In particular, $C_{\Gamma,\Gamma_{0}}$ is a complete CAT(0) space as a subspace of $C_{\Gamma}$.

### v3

@@ 2.2. Defining graphs of Artin groups
We now have a finite simple labeled graph $\Gamma$ with the vertex set $V(\Gamma)=V$, the edge set $E(\Gamma)=\bigsqcup_{m\in\mathbb{Z}_{\geq}2}\tilde{\mu}^{-1}(m)$, and the labeling $\mu:=\tilde{\mu}|_{E(\Gamma)}$. The Artin group $A_{\Gamma}$ associated with $\Gamma$ is defined by the presentation (1.1), and $\Gamma$ is called the defining graph of $A_{\Gamma}$.
For convenience, we define two other finite simple graphs, $\Gamma^{c}$ and $\Gamma^{t}$, as follows. $\Gamma^{c}$ is the finite simple graph with the vertex set $V(\Gamma^{c})=V$ and the edge set $E(\Gamma^{c})=\tilde{\mu}^{-1}(\infty)$. This is called the complement graph of $\Gamma$. The Coxeter graph $\Gamma^{t}$ is the finite simple graph with the vertex set $V(\Gamma^{t})=V$ and the edge set $E(\Gamma^{t})=\bigsqcup_{m\in\mathbb{Z}_{\geq 3}\cup\{\infty\}}\tilde{\mu}^{-1}(m)$.
@@ 2.3. Properties of Artin groups
We summarize some properties of Artin groups. Let $A_{\Gamma}$ be an Artin group associated with a defining graph $\Gamma$ with the vertex set $V(\Gamma)=V$ and the edge set $E(\Gamma)=E$, as in Section 2.2. Lemma 2.11 is essentially used in the proof of Theorem 1.2.
By the work of van der Lek (see [35], [41]), the following are well known and often used without mention in this paper.
@@
    For any subset $U\subset V$, the subgroup of $A_{\Gamma}$ generated by $U$, called a standard parabolic subgroup, is itself an Artin group associated with the full subgraph of $\Gamma$ spanned by $U$. (We denote this subgroup by $A_{U}$. When $U$ is empty, we define $A_{\emptyset}=\{1\}$.)
@@
Take an element $v\in V$ and a subset $U\subset V$. Suppose that $v\not\in U$. Set $W:=\{w\in U\mid\mu(v,w)=2\}$. Then $A_{U}\cap vA_{U}v^{-1}=A_{W}$. In particular, for any $u\in U\setminus W$, $A_{U}\cap vA_{U}v^{-1}\subset A_{U\setminus\{u\}}$.
@@
Clearly we have $A_{W}\subset A_{U}\cap vA_{U}v^{-1}$. In terms of [19], $v$ is $(U,U)$-reduced in the Coxeter group $W_{\Gamma}$. Applying [19, Proposition 3.3] to our setting, we have some $U^{\prime}\subset U$ such that $vA_{U}v^{-1}\cap A_{U}=A_{U^{\prime}}$. Hence $W\subset U^{\prime}$. Assume that $W\subsetneq U^{\prime}$. Take $u^{\prime}\in U^{\prime}\setminus W$. Then $u^{\prime}\in A_{U^{\prime}}=vA_{U}v^{-1}\cap A_{U}$ and thus $v^{-1}u^{\prime}v\in A_{U}$. By $v^{-1}u^{\prime}v\in A_{\{v,u^{\prime}\}}$ and $\{v,u^{\prime}\}\cap U=\{u^{\prime}\}$, we have $v^{-1}u^{\prime}v\in A_{\{v,u^{\prime}\}}\cap A_{U}=A_{\{u^{\prime}\}}$. Therefore, we have $m\in\mathbb{Z}$ such that $v^{-1}u^{\prime}v=(u^{\prime})^{m}$. Since the surjection $A_{\{v,u^{\prime}\}}\to A_{\{u^{\prime}\}}$ defined by $v,u^{\prime}\mapsto u^{\prime}$ gives $v^{-1}u^{\prime}v\mapsto u^{\prime}$ and $(u^{\prime})^{m}\mapsto(u^{\prime})^{m}$, we see $u^{\prime}={(u^{\prime})}^{m}$ and $m=1$. Hence we have $v^{-1}u^{\prime}v=u^{\prime}$ in $A_{\Gamma}$. This means $\mu(v,u^{\prime})=2$, contradicting $u^{\prime}\notin W$. Hence $U^{\prime}=W$. Therefore, $vA_{U}v^{-1}\cap A_{U}=A_{W}$. ∎
@@ 2.4. Clique-cube complexes and actions on them
The clique-cube complex $C_{\Gamma}$ is the cube complex whose vertices (i.e., 00-dimensional cubes) are cosets $gA_{U}$ ($g\in A_{\Gamma},U\in\Delta_{\Gamma}$), where two vertices $gA_{U}$ and $hA_{U^{\prime}}$ are joined by an edge (i.e., a $1$-dimensional cube) in $C_{\Gamma}$ if $gA_{U}\subset hA_{U^{\prime}}$ and $U$ and $U^{\prime}$ differ by a single generator. Note that, in this case, we can always replace $h$ by $g$ so that $hA_{U^{\prime}}=gA_{U^{\prime}}$. More generally, two vertices $gA_{U}$ and $gA_{U^{\prime}}$ with $gA_{U}\subset gA_{U^{\prime}}$ span a #$\#(U^{\prime}\setminus U)$-dimensional cube $[gA_{U},gA_{U^{\prime}}]$ in $C_{\Gamma}$.
See Figure 3.
Figure 3. For the above graph $\Gamma$, the below cube complex is a part of $C_{\Gamma}$ that is a fundamental domain of $C_{\Gamma}$ with respect to the action of $A_{\Gamma}$. We define a subcomplex $C_{\Gamma,\Gamma_{0}}$ of $C_{\Gamma}$ in Section 3. The subcomplex consisting of colored vertices and thick edges is a part of $C_{\Gamma,\Gamma_{0}}$, which is a fundamental domain of $C_{\Gamma,\Gamma_{0}}$ with respect to the action of $A_{\Gamma}$.
We endow a metric on $C_{\Gamma}$ in the standard way for a cube complex. That is, we identify all cubes in $C_{\Gamma}$ with Euclidean unit cubes and give $C_{\Gamma}$ the path metric. Note that $C_{\Gamma}$ is connected.
Theorem 2.13 ([14, Theorem 2.2]).
The clique-cube complex $C_{\Gamma}$ is a finite dimensional CAT(0) cube complex. In particular, $C_{\Gamma}$ is a complete CAT(0) space.
The group $A_{\Gamma}$ acts on the clique-cube complex $C_{\Gamma}$ by left multiplication, $h\cdot gA_{U}=(hg)A_{U}$. This action preserves the cubical structure and is isometric. The action is also cocompact with a fundamental domain $\bigcup_{U\in\Delta_{\Gamma}}[A_{\emptyset},A_{U}]$, where $[A_{\emptyset},A_{U}]$ is a #$\#U$-dimensional cube spanned by the two vertices $A_{\emptyset}$ and $A_{U}$ in $C_{\Gamma}$. See Figure 3. However, the action is not proper. In fact, the stabilizer of a vertex $gA_{U}$ is the conjugate subgroup $gA_{U}g^{-1}$, so all vertices other than those in the orbit of $A_{\emptyset}$ have infinite stabilizers. We also note that $C_{\Gamma}$ is not a proper metric space because it contains vertices of infinite valence. Additionally, $C_{\Gamma}$ has infinite diameter if and only if $\Gamma$ itself is not a clique.
@@
Each edge in $C_{\Gamma}$ can be labeled with a generator in $V$. Indeed, the edge between $gA_{U}$ and $gA_{U\sqcup\{v\}}$ is labeled by $v$. Any two parallel edges in a cube have the same label, so we can also label the hyperplane dual to such an edge by $v$ and say that such a hyperplane is of $v$-type. Every hyperplane of $v$-type is a translation of the hyperplane dual to the edge between $A_{\emptyset}$ and $A_{\{v\}}$. If a hyperplane of $v$-type crosses another hyperplane of $v^{\prime}$-type, then $(v,v^{\prime})\in E$. In particular, two different hyperplanes of the same type do not cross each other.
@@ 2.5. Joins of graphs
A simple graph $\Gamma$ is said to be decomposable as a join if there exists an index set $\Lambda$ with #$\#\Lambda\geq 2$ and subgraphs $\Gamma_{\alpha}$ ($\alpha\in\Lambda$) of $\Gamma$ such that $\Gamma=\ast_{\alpha\in\Lambda}\Gamma_{\alpha}$. This is called a join decomposition of $\Gamma$ into factors $\Gamma_{\alpha}$ ($\alpha\in\Lambda$). $\Gamma$ is said to be indecomposable as a join if it is not decomposable as a join.
@@
A simple graph $\Gamma$ is indecomposable as a join if and only if its complement graph $\Gamma^{c}$ is connected.
@@
For convenience, even if $\Gamma$ is indecomposable as a join, we take $\Lambda$ consisting of a single index $\alpha$, define $\Gamma_{\alpha}=\Gamma$ and regard $\Gamma=\ast_{\alpha\in\Lambda}\Gamma_{\alpha}$ as the unique join decomposition of $\Gamma$ (by a single indecomposable factor).
Figure 4. The defining graphs $\Gamma$ and the associated Coxeter graphs $\Gamma^{t}$. Each $\Gamma$ has a join decomposition $\Gamma=\Gamma_{0}\ast\Gamma_{1}$, where $\Gamma_{0}$ is the maximal clique factor and $\Gamma_{1}$ is an indecomposable factor with two vertices.
Figure 5. The defining graphs $\Gamma$ and the associated Coxeter graphs $\Gamma^{t}$. Each $\Gamma$ has a join decomposition $\Gamma=\Gamma_{0}\ast\Gamma_{1}\ast\Gamma_{2}$, where $\Gamma_{0}$ is the maximal clique factor and each of $\Gamma_{1}$ and $\Gamma_{2}$ is an indecomposable factor with two vertices.
@@
Let $\Gamma$ be a simple graph. For the unique join decomposition into indecomposable factors $\Gamma=\ast_{\alpha\in\Lambda}\Gamma_{\alpha}$, we set $\Lambda_{0}:=\{\alpha\in\Lambda\mid\ \#\Gamma_{\alpha}=1\}$ and $\Gamma_{0}:=\ast_{\alpha\in\Lambda_{0}}\Gamma_{\alpha}$. We call $\Gamma_{0}$ the maximal clique factor of $\Gamma$. Also we set $\Lambda_{\ast}:=\Lambda\setminus\Lambda_{0}$ and $\Gamma_{\ast}:=\ast_{\alpha\in\Lambda_{\ast}}\Gamma_{\alpha}$. We call $\Gamma$ a cone if $\Lambda_{0}\neq\emptyset$.
See Figures 1, 2, 4 and 5.
We have $\Gamma=\Gamma_{0}\ast\Gamma_{\ast}$. Clearly, $\Gamma$ is a clique if and only if $\Lambda=\Lambda_{0}$.
2.6. Notation on matrices
For natural numbers $k$, $m$, and a set $X$, let $\mathcal{M}(k,m;X)$ be the set of all $k\times m$ matrices whose entries are in $X$. For $i\in\{1,\ldots,k\}$ and $l\in\{1,\ldots,m\}$, let $M[i,l]$ be the $(i,l)$-component of $M$. Let $M[i,\ ]$ be the $i$-th row and $M[\ ,l]$ be the $l$-th column of the matrix $M$. For $l_{1},l_{2}\in\{1,\ldots,m\}$ with $l_{1}\leq l_{2}$, we define $M[\ ,l_{1}:l_{2}]\in\mathcal{M}(k,l_{2}-l_{1}+1;X)$ by
(2.1) $\left(M[\ ,l_{1}:l_{2}]\right)[\ ,l]=M[\ ,l_{1}-1+l]\quad(1\leq l\leq l_{2}-l_{1}+1).$
For two matrices $M_{j}\in\mathcal{M}(k,m_{j};X)$ $(j=1,2)$ with the same number of rows, we define $cbind(M_{1},M_{2})\in\mathcal{M}(k,m_{1}+m_{2};X)$ as a matrix constructed from $M_{1},M_{2}$ by binding the columns. That is,
(2.2) $cbind(M_{1},M_{2})[\ ,l]=\begin{cases}M_{1}[\ ,l]\quad(1\leq l\leq m_{1})\\ M_{2}[\ ,l-m_{1}]\quad(m_{1}+1\leq l\leq m_{1}+m_{2}).\end{cases}$
For matrices $M_{j}\in\mathcal{M}(k,m_{j};X)$ $(1\leq j\leq r)$ with the same number of rows, we define $cbind(M_{1},\ldots,M_{r})\in\mathcal{M}(k,m_{1}+\cdots+m_{r};X)$ inductively by
(2.3) $cbind(M_{1},\ldots,M_{r})=cbind\left(cbind(M_{1},\ldots,M_{r-1}),M_{r}\right).$
Let $A$ be a group and let $M\in\mathcal{M}(k,m;A)$. For any subset $B\subset A$, which is typically a subgroup or a generating set, we naturally regard $\mathcal{M}(k,m;B)\subset\mathcal{M}(k,m;A)$. For two matrices $M_{1},M_{2}\in\mathcal{M}(k,m;A)$ of the same size, $M_{1}\odot M_{2}\in\mathcal{M}(k,m;A)$ denotes the Hadamard product of $M_{1}$ and $M_{2}$. That is,
(2.4) $(M_{1}\odot M_{2})[i,l]=M_{1}[i,l]\cdot M_{2}[i,l]\in A.$
We define a map $prod:\mathcal{M}(k,m;A)\to A$ inductively as follows. When $m=1$, let
(2.5) $prod(M)=M[1,1]\cdot M[2,1]\cdot\cdots\cdot M[k,1]\in A.$
If $prod$ is defined on $\mathcal{M}(k,m-1;A)$, we define $prod:\mathcal{M}(k,m;A)\to A$ by
(2.6) $prod(M)=prod(M[\ ,1:m-1])\cdot prod(M[\ ,m])\in A.$
@@ 3. Reduced clique-cube complexes and actions on them
In this section, we introduce and study reduced clique-cube complexes. Let $A_{\Gamma}$ be an Artin group associated with a defining graph $\Gamma$ with the vertex set $V(\Gamma)=V$ and the edge set $E(\Gamma)=E$, as in Section 2.2. Let $\Gamma_{0}$ be the maximal clique factor of $\Gamma$. We consider the join decomposition $\Gamma=\Gamma_{0}\ast\Gamma_{\ast}$. We set $V_{0}=V(\Gamma_{0})$.
@@
$\Delta_{\Gamma,\Gamma_{0}}=\{U\subset V\mid U\in\Delta_{\Gamma},\ U\supset V_{0}\}.$
Let $C_{\Gamma,\Gamma_{0}}$ be the subcomplex of $C_{\Gamma}$ spanned by vertices $gA_{U}$ with $g\in A_{\Gamma}$ and $U\in\Delta_{\Gamma,\Gamma_{0}}$. We call $C_{\Gamma,\Gamma_{0}}$ the reduced clique-cube complex.
See Figure 3.
When $\Gamma$ is not a cone; that is, $\Gamma=\Gamma_{\ast}$, we have $C_{\Gamma,\Gamma_{0}}=C_{\Gamma}$. When $\Gamma$ is a clique; that is, $\Gamma=\Gamma_{0}$, we have $C_{\Gamma,\Gamma_{0}}=C_{\Gamma,\Gamma}=\{A_{\Gamma}\}$. In this paper, we are mainly interested in the case where $\Gamma$ is not a clique.
We note that the action of $A_{\Gamma}$ on $C_{\Gamma}$ preserves $C_{\Gamma,\Gamma_{0}}$. We consider the induced action of $A_{\Gamma}$ on $C_{\Gamma,\Gamma_{0}}$. Then $\bigcup_{U\in\Delta_{\Gamma,\Gamma_{0}}}[A_{V_{0}},A_{U}]$ is a fundamental domain of $C_{\Gamma,\Gamma_{0}}$ with respect to the action of $A_{\Gamma}$.
@@
If $\Gamma$ is a clique, then $C_{\Gamma,\Gamma_{0}}=\{A_{\Gamma}\}$ is clearly connected. Suppose that $\Gamma$ is not a clique. Let $U\in\Delta_{\Gamma,\Gamma_{0}}$. Let $g\in A_{\Gamma}$. First, we show that there exists a path in $C_{\Gamma,\Gamma_{0}}$ from $A_{V_{0}}$ to $gA_{V_{0}}$. Suppose that $g\notin A_{V_{0}}$; otherwise, we would have $A_{V_{0}}=gA_{V_{0}}$. Let $g=v_{1}^{\varepsilon_{1}}\cdots v_{n}^{\varepsilon_{n}}$, where $n\in\mathbb{N}$, $v_{i}\in V$, $\varepsilon_{i}\in\{1,-1\}$ for every $1\leq i\leq n$. By induction on $n$, we suppose that there exists a path in $C_{\Gamma,\Gamma_{0}}$ from $A_{V_{0}}$ to $v_{1}^{\varepsilon_{1}}\cdots v_{n-1}^{\varepsilon_{n-1}}A_{V_{0}}$. If $v_{n}\in V_{0}$, then
$v_{1}^{\varepsilon_{1}}\cdots v_{n-1}^{\varepsilon_{n-1}}A_{V_{0}}=v_{1}^{\varepsilon_{1}}\cdots v_{n}^{\varepsilon_{n}}A_{V_{0}},$
and the induction proceeds. If $v_{n}\notin V_{0}$, since $V_{0}\sqcup\{v_{n}\}\in\Delta_{\Gamma,\Gamma_{0}}$, there exists a path
$\left(v_{1}^{\varepsilon_{1}}\cdots v_{n-1}^{\varepsilon_{n-1}}A_{V_{0}},\ v_{1}^{\varepsilon_{1}}\cdots v_{n-1}^{\varepsilon_{n-1}}A_{V_{0}\sqcup\{v_{n}\}}=gA_{V_{0}\sqcup\{v_{n}\}},\ gA_{V_{0}}\right)$
in $C_{\Gamma,\Gamma_{0}}$ from $v_{1}^{\varepsilon_{1}}\cdots v_{n-1}^{\varepsilon_{n-1}}A_{V_{0}}$ to $gA_{V_{0}}$. By concatenating these two paths, we obtain a path from $A_{V_{0}}$ to $gA_{V_{0}}$ in $C_{\Gamma,\Gamma_{0}}$. Next, we show the existence of a path in $C_{\Gamma,\Gamma_{0}}$ from $gA_{V_{0}}$ to $gA_{U}$. Let $U=V_{0}\sqcup\{u_{1},\ldots,u_{m}\}$ where $u_{i}\neq u_{j}$ for $1\leq i\neq j\leq m$. Then there exists a path
$\left(gA_{V_{0}},gA_{V_{0}\sqcup\{u_{1}\}},gA_{V_{0}\sqcup\{u_{1},u_{2}\}},\ldots,gA_{U}\right)$
in $C_{\Gamma,\Gamma_{0}}$ from $gA_{V_{0}}$ to $gA_{U}$. Therefore, it follows that $C_{\Gamma,\Gamma_{0}}$ is a connected subcomplex of $C_{\Gamma}$. ∎
We endow $C_{\Gamma,\Gamma_{0}}$ with a metric in the standard way for a cube complex. That is, we identify all cubes in $C_{\Gamma,\Gamma_{0}}$ with Euclidean unit cubes and give $C_{\Gamma,\Gamma_{0}}$ the path metric. Note that $C_{\Gamma,\Gamma_{0}}$ is connected.
@@
Let $X$ be a geodesic space and $W$ be a subset of $X$. $W$ is convex in $X$ if for every $w_{1},w_{2}\in W$, every geodesic from $w_{1}$ to $w_{2}$ is contained in $W$.
@@
$C_{\Gamma,\Gamma_{0}}$ is convex in $C_{\Gamma}$. In particular, $C_{\Gamma,\Gamma_{0}}$ is a finite dimensional CAT(0) cube complex and a complete CAT(0) space.